OpenStudy (anonymous):

solve the radical equations: √3x + 10 = 2+√x+2

OpenStudy (jiteshmeghwal9):

Is this your question? \(\sqrt{3}x+10=2+\sqrt{1}x+2\)

OpenStudy (jiteshmeghwal9):

@glares25 ?

OpenStudy (anonymous):

yes

OpenStudy (amorfide):

you have square roots so you should square both sides then move all constants onto one side and all unknowns on the other side then work from there

OpenStudy (jiteshmeghwal9):

Take the radicals at one side & constants at one side & then square both sides & then solve :)

hartnn (hartnn):

are u sure its not this \(\sqrt{3x+10}=2+\sqrt{x+2}\) and its actually what jitesh wrote ?

OpenStudy (anonymous):

@hartnn yes that one

hartnn (hartnn):

then u directly square both sides, what u get ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok, thank you so much .. i can work on my own ,thanks :)

hartnn (hartnn):

sure? coz we haven't even started yet....